### v1

# WakenLLM: A Fine-Grained Benchmark for Evaluating LLM Reasoning Potential and Reasoning Process Stability

Abstract
Large Language Models (LLMs) frequently output the label Unknown, yet current evaluations focus almost exclusively on whether such answers are honest rather than why they arise. This blurs two distinct cases: (i) an input that is genuinely indeterminate and (ii) a solvable problem that the model fails to resolve. We call this phenomenon Vague Perception. And thus we introduce a framework that quantifies the proportion of Unknown responses attributable to model incapacity and tests whether guided stimulation can convert them into either correct (Known) or intrinsically indeterminate outcomes. By separating these sources of uncertainty, our method provides a clearer picture of LLM reasoning limits and their potential for improvement. As we get a theoretical accuracy of reasoning task on different LLMs, we apply different methods to test whether the model can reach the accuracy given a baseline framework. Our work is meaningful in exploring the true reasoning ability of LLMs and providing a new perspective on solving the Vague Perception phenomenon.
Introduction
Large Language Models (LLMs) such as GPT-3, PaLM, and GPT-4 have made remarkable progress across wide range of language understanding and generation tasks, exhibiting impressive zero-shot and few-shot reasoning skills on benchmarks like BIG-Bench, MMLU, and GSM8K (Brown et al. 2020; OpenAI 2023; Chowdhery et al. 2022; Srivastava and et al. 2022; Hendrycks and et al. 2021; Cobbe et al. 2021). These capabilities are further enhanced by prompting techniques—including Chain-of-Thought, Self-Consistency, and Program-Aided Reasoning—that explicitly elicit intermediate reasoning steps (Kojima et al. 2022; Wei et al. 2022; Wang et al. 2023; Lewkowycz and et al. 2022). Despite this progress, most evaluations still treat the model as a black box, asking only whether the final answer is correct, truthful, or aligned with human intent (Mitchell et al. 2022; Jiang et al. 2024; Bubeck 2023).
Fine-grained reasoning reliability remains under-explored. Empirical studies have shown that LLMs frequently hallucinate, hedge, or output “Unknown” even when the context contains sufficient evidence, while at other times they confidently answer questions that are fundamentally under-specified (Lin 2023; Ye and et al. 2023). This paper formalizes this phenomenon as Vague Perception: a mismatch between a model’s internal reasoning process and the logical sufficiency of the input. Therefore, a critical gap lies in understanding how LLMs parse complex information, decide that evidence is (in)sufficient, and transition from partial inference to a definitive judgment when faced with ambiguity or uncertainty.
Figure 1: Definition of Vague Perception: when an LLM misjudges a Verifiable sample as (“Unknown”), or it is unable to tell why a sample should be categorized as unknown inputting Unverifiable Samples. The figure presents two representative examples attributed to distinct causes: Fact Understanding and Reasoning Gap—highlighting 1. Limited ability to solve a problem. 2.Misunderstandings of facts versus failures to link them.
In this work, we analyze the root causes of deep-reasoning failures behind Vague Perception and propose fine-grained mitigation strategies. To enable systematic measurement, we introduce a new benchmark that spans four subject domains and six subcategories, offered in two complementary formats: (i) Fact-based benchmark, where relevant information is listed as discrete atomic statements (e.g., “A is X, B is Y”); and (ii) Story-based benchmark, where the same facts are embedded in a coherent passage. This design lets us probe how presentation style and information density influence the emergence and eventual resolution of Vague Perception.
Paradigm Guidance Questioning Stimulation Dynamic Samples Selection Reflection Stage-wise Examining Fine-Grained Decomposing
CoT (Wei et al. 2022)
ToT (Yao et al. 2023)
Faithful-CoT (Lyu et al. 2023)
Self-Refine (Madaan et al. 2023)
Reflexion (Shinn et al. 2023)
Decomposed Prompting (Khot et al. 2023)
Active Prompt (Diao et al. 2023)
SCORE (Zhang et al. 2024)
ProCo (Wu et al. 2024)
CoTAL (Cohn et al. 2025)
Cochain (Zhao et al. 2025)
ISP2 (Zhu et al. 2025)
[…]
Table 1: Comparison of different LLM guidance reasoning frameworks.
We construct benchmarks in which half of the samples has T/F labels and the other half is Unknown. These samples are then fed into the model to extract the corresponding Vague Perception samples. Subsequently, the selected instances were processed via a two-stage stimulation mechanism aimed at “waking up” the model’s latent reasoning abilities: In stage 1, we enumerate explicit rules that the LLM should follow in order to reanalyze the problems.
By measuring Awaken Ability, we present a novel perspective on evaluating LLM reasoning, one that considers not only the model’s initial accuracy, but also its potential to be guided from error to correctness.
We propose WakenLLM as a strong baseline to explore the reasoning abilities of LLMs among Vague Perception samples, yielding comparable performance with GPT-3.5. Furthermore, we evaluate the performance of over 30 LLMs across these reasoning methods on TableBench, highlighting that both open-source and proprietary LLMs require substantial improvements to meet real-world demands. Notably, even one of the most advanced models, GPT-4o, achieves only a modest score.
To summarize, our contributions are threefold:
  • We perform a comprehensive investigation of Vague Perception on both Fact-based and Story-based benchmarks, analyzing its influence on reasoning in casual and professional benchmarks.
  • We introduce WakenLLM, a unified framework to evaluate the ”Awaken Ability” of LLMs—their capacity to produce correct answers after stimulation—representing a principled approach to handling Vague Perception samples.
  • Through a systematic analysis of models with varying parameter scales, we establish benchmark metrics that more precisely quantify the challenges posed by Vague Perception. The resulting insights not only improve reasoning accuracy but also advance our understanding of LLM trustworthiness and in-depth reasoning paradigm.
Evaluation Framework
The WakenLLM evaluation framework, illustrated in Figure 2, is composed of two successive stages. It evaluates the model’s Awakening Ability and systematically addresses the challenge of Vague Perception. Specifically, the first stage assesses the Overall Benchmark Converting Rate, while the second investigates the impact of stimulation under the Remind-then-Guide (RtG) paradigm on both label and reasoning process conformity.
Overall Benchmark Converting Rate
Our framework begins with a model-specific pre-processing stage that identifies Vague Perception samples—cases where the LLM initially mislabels (e.g., saying Unknown when the sample is verifiable, or failing to explain an unverifiable case). Since this process is performed separately for each model, the set of Vague Perception samples differs across models. The converting results fall into three categories: (1) The model fails to generate a coherent logic chain, which we call Unexcited Converting (UC). (2) The model produces a false logic chain and an incorrect final answer, denoted as False Converting (FC). (3) The model outputs a correct label under stimulation, called True Converting (TC). To quantify these proportions, we use UCR, FCR, and TCR to represent the rates of each type of conversion. Given our focus on the genuine reasoning correction of the model, we elaborate only on TCR. UCR and FCR follow analogous logic. In order to measure the ability of judging both verifiable and unverifiable data, we denote 2 separate metrics for each, defining True Converting Rate(TCR) in N-th stage as follows:
$\begin{gathered}TC_{f}^{n}\;=\;\sum_{y_{f}\in D_{\mathrm{VP}}^{n}}\mathbf{1}\bigl{(}\hat{y}_{f}^{n}=y_{f}\bigr{)},\\[6.0pt] TCR_{f}^{n}\;=\;\frac{\bigl{|}TC_{f}^{n}\bigr{|}}{\bigl{|}D_{\mathrm{VP}}^{n}\bigr{|}}\,,\\[6.0pt] \text{where}\quad f\in\{u,v\},\quad n\in\{1,2\}.\end{gathered}$
Figure 2: Overview of the proposed pipeline. We begin with a balanced dataset with Fact and Story-based dataset, sampling verifiable (T/F) and unverifiable (Unknown) instances at a 1 : 1 ratio. The pipeline then (1) detects cases that expose the model’s Vague Perception during reasoning and (2) applies a two-stage remedy: Stage 1 uses stimulation to re-evaluate the mislabeled samples, while Stage 2 gathers the remaining False Converting samples from Stage 1 for model self-reflection. Stimulation Types design vary with data form.
Here, $TC_{v}$ quantifies input samples that have been correctly classified, while $TC_{u}$ refers to unverifiable samples where an Unknown label has been correctly assigned. $N$ denotes the stimulation stage, which can be either one or two. $D_{\text{VP}}^{(n)}$ represents the input of Vague Perception samples in the $n$-th stage. We can tell that $D_{\mathrm{VP}}^{2}=FC^{2}$.
It should be noted that both $TCR_{v}$ and $TCR_{u}$ quantities represent the results at the current stage, rather than the cumulative sum across both stages. To integrates performance across both stages, we define Overall Converting Rate (OCR) as follows:
$OCR=\frac{\lvert TC^{1}\cup TC^{2}\rvert}{\lvert D_{\mathrm{vp}}\rvert},\quad TC_{1}\cap TC_{2}=\emptyset.$
It refers to the portion of under-stimulated samples turn out to be predicted right. By doing this, we can obtain the theoretical potential accuracy output by a model.
Remind-then-Guide (RtG) Conformity Stimulating
Label Offering
We test whether Vague Perception samples display a better or worse stability—i.e. whether they are more easily misled—relative to ordinary samples. According to previous work (Weng, Chen, and Wang 2025), all models demonstrate Conformity. In this part, we first inform the model that it has misunderstood the sample, and then provide a randomly assigned label with varying probabilities, to test whether this misdirection affects its final decision.
We evaluate this phenomenon by setting up two aspects of one metrics:
X can be v or u, indicating whether a sample is objectively verifiable or unverifiable on its own. Here, $S$ denotes the model is under the RTG stimulation.
Reasoning Process (RP) Offering
We examine how effectively the reasoning chain generated in Stage 1 serves as inspiration or guidance for Stage 2 when it is supplied as additional input. To ensure that the model demonstrates genuine understanding rather than simply being prompted to output the correct answer by accident, we propose two metrics.
Correct Guiding Rate (CGR): This metric assesses whether the reasoning process output in Stage 1 can assist Stage 2 in stimulating samples to better avoid prior mistakes.
Reasoning Process Coherence (RPC): This metric involves adding the reasoning process from Stage 1 to re-run Stage 1, measuring the alignment of the reasoning process provided by Stage 1 to assess its stability. where R1 denotes the input contains the reasoning process output by Stage 1.
Experiments
Benchmark
To explore the effects of the Vague Perception phenomenon on different types of logical data, we classify them based on their labels: (1) Verifiable Data, where the samples have a clear True/False label indicating the validity of the reasoning; and (2) Unverifiable Data, where the label is Unknown and samples are inherently unsolvable.
Furthermore, as shown in Figure 3, to better understand how large language models (LLMs) reason across different logical forms, we categorize the whole benchmarks as follows:
Fact-based Benchmark.
FLD (Morishita et al. 2024) supplies three kinds of sample labels annotated as True, False, or Unknown. We randomly sample 600 items: 300 verifiable (True/False) and 300 marked Unknown due to insufficient evidence. FOLIO (Han et al. 2022) is a factual benchmark with the same 3-way labeling. Following an identical protocol, we draw 640 instances—320 verifiable claims and 320 Unknown.
Figure 3: Overall benchmark distribution. Our dataset contains 2,840 samples grouped into two top-level task types: story-based (56.3%) and fact-based (43.7%). A detailed breakdown of the “Else” category is provided in the Appendix.
Story-based Benchmark.
For ScienceQA (Lu et al. 2022) we construct two balanced splits to capture both Science and Arts reasoning. The Science split contains 400 Physics and 400 Biology questions; the Arts split holds 400 Figurative-Language and 400 Writing-Strategies questions. Within each split exactly half of the items are answerable from the passage, while the remaining half are deliberately labelled Unknown, mirroring the fact-verification setting and enabling a unified evaluation protocol.
Our experimental design considers the differing characteristics of Fact-based benchmark and Story-based benchmark. For fact-based tasks, which models tend to process the entire context holistically, so providing additional stimulation can enhance performance. In contrast, for story-based tasks—often rooted in professional knowledge domains and relying on models’ inherent knowledge—we minimize guidance to avoid interfering with the model’s independent reasoning process.
Experimental Setup
Our evaluation covers both closed-source (ChatGPT-3.5-turbo and GPT-4o) and open-source models (Llama-3.1, Gemma-2, and Qwen-2.5 series) (OpenAI 2022; Touvron and Team 2025; Google DeepMind 2025; Bai and Team 2025). Two stimulating strategies are implemented:
  • Concise Stimulation: Basic error notification without correction guidance.
  • Detailed Stimulation: Enhanced detailed approach featuring:
    • Listed Conditions Guidance
    • Structured Reasoning Constraints
    • Conditional Verification Steps
The experimental protocol differs between approaches. Concise Stimulation uses a similar format for two-stage error notification process, while Detailed Stimulation employs a different two-stage workflow: a standard prediction followed by an interrogative refinement. The latter stimulates analytical reasoning through rhetorical questioning and step-by-step error analysis, creating a clear methodological contrast.
Results on Overall Benchmark Converting Rate
As shown in Table 3, we report the overall benchmark Converting Rate, which means 4 metrics in three systems: Unexcited Converting Rate (UCR), False Converting Rate (FCR), and True Converting Rate for both verifiable and unverifiable data (TCRv, TCRu).
Benchmark Samples Target Form Writing Type
FLD 600 Conclusion Abstract Generated
FOLIO 640 Conclusion Plain Real
Science 800 Q&A Plain Real
Arts 800 Q&A Plain Real
Table 2: Basic info of each benchmark.
Model Dataset Vanilla RtG (Label) RtG (RP)
TCR1 TCR2 OCR Confv Confu CGR RPC
GPT-3.5 FLD 12.61 14.43 25.22 64.25 3.33 -12.30 10.84
FOLIO 14.41 42.10 50.44 16.81 31.97 -42.10 -6.30
Science 21.05 3.33 21.63 19.22 1.96 -3.33 0.59
Arts 26.04 4.00 27.10 14.24 -3.88 0.00 3.12
GPT-4o FLD 18.88 28.19 41.65 8.17 37.38 -23.50 -6.48
FOLIO 37.34 44.40 65.16 9.63 51.90 -44.77 0.00
Science 24.73 7.76 28.96 11.74 4.58 -7.16 -3.68
Arts 18.60 8.24 24.17 8.07 23.31 -8.24 1.14
LLaMA-3.1-8B FLD 41.97 26.03 48.90 50.04 11.04 -18.56 1.09
FOLIO 45.86 11.40 50.84 38.73 13.61 -3.07 5.52
Science 45.94 5.95 50.61 38.60 51.44 -5.95 12.91
Arts 24.26 8.57 25.75 30.99 9.02 -4.77 11.30
LLaMA-3.1-405B FLD 20.65 53.73 41.50 9.35 57.63 -48.28 -17.66
FOLIO 28.76 29.92 49.22 6.11 23.01 -17.69 0.00
Science 32.85 2.50 33.03 9.02 18.50 0.19 0.71
Arts 4.40 0.00 4.40 0.81 28.91 0 -0.77
Qwen-2.5-7B FLD 48.31 40.00 68.53 22.06 42.34 -31.67 -5.61
FOLIO 19.73 36.84 47.36 4.92 38.71 -36.84 14.48
Science 15.83 10.00 24.08 55.21 0.00 -2.98 14.59
Arts 29.34 25.37 36.44 43.39 7.20 -13.27 8.47
Qwen-2.5-72B FLD 48.20 9.80 51.21 33.08 9.14 -9.80 1.21
FOLIO 31.51 11.37 38.36 10.65 2.20 -6.82 -1.37
Science 10.81 3.12 11.20 33.42 22.07 0.00 -0.38
Arts 25.00 9.76 25.56 12.52 24.12 0.00 -1.95
Table 3: Comparison of six LLMs across four task subclasses (FLD, FOLIO, Science, Arts), grouped by evaluation settings.
Results on RtG Conformity Stimulating
As illustrated before, RtG module contains two parts: We Remind the model that it mis-labeled the samples before, and we apply conformity test thereafter.
Label Conformity
Judging from the table, we observe that even when LLMs are reminded of their previous incorrect answers, they still demonstrate a strong tendency toward conformity. Moreover, distinct conformity patterns emerge between verifiable and unverifiable samples. One key influencing factor appears to be the model size: models with smaller parameter sizes, such as GPT-3.5 and LLaMA-3.1-8B, are more prone to being misled on verifiable samples involving T/F label reasoning. In contrast, models with larger parameter sizes, such as GPT-4o and LLaMA-3.1-405B, tend to be more easily misled when the label is Unknown. Meanwhile, Qwen series shows a distinctive feature different from the above two kinds, we infer it is because parameters gap between Qwen is much smaller than those.
Reasoning Process Conformity
When we input the previous reasoning process and explicitly ask the LLM to reflect on its mistakes, the resulting CGR values are mostly below zero. This indicates that even after being reminded of a past error and provided with its own flawed reasoning, the LLM still tends to repeat the same incorrect conclusion, leading to a noticeable decline in accuracy.
Furthermore, we investigate whether the reasoning process generated by LLMs is generally stable. Our results reveal that the reasoning process is often highly unstable. Changes in accuracy can largely be attributed to differences in how the problem is interpreted. Notably, for unverifiable samples, the reasoning process tends to be unreliable — feeding it back into the model typically leads to a further drop in accuracy. This exposes a hidden issue: when LLMs automatically generate solutions using a reasoning process, unverifiable problems are particularly susceptible to unreliability. Even for verifiable samples, the reasoning may remain unstable.
Discussion
As shown in the line chart of Fig 4, by comparing Stage 1 and 2 True Converting Rate (TCR1 and TCR2) across different benchmark types, we observe clear divergences: Fact-based benchmarks tend to achieve a markedly higher TCR2, implying that a great amount of mis-predicted samples got predicted right in Stage 2, where the model can effectively refine the outputs produced in Stage 1. In contrast, Story-based benchmarks’ converting rate excel in Stage 1 but show a substantially lower TCR2. Thus, distinct stimulation strategies are recommended: For fact-based benchmarks, broader exploratory freedom is acceptable in Stage 1. “Making mistakes at first is allowed”, since later refinement proves beneficial; For Story-based benchmarks, the model must “do it right the first time,” as the improvement of subsequent stage is limited. Hence, optimization priorities diverge: Fact-based tasks profit from a two-stage pipeline that revises early missteps, whereas story-based tasks demand maximizing single-pass accuracy.
Detailed Analysis
Benchmark-wise Vague Perception
As each model differs in its reasoning ability, we first quantify the percentage of data models induce Vague Perception. Fig 4 reports the percentage for GPT-4o—the most widely used LLM at present—while complete results for the remaining models appear in the appendix.
First, Vague Perception is prevalent across all benchmark. FOLIO exhibits a relatively lower percentage. In the FLD and Arts benchmarks—which focuses purely on language Understanding, the model demonstrates greater uncertainty regarding explaining why certain samples are labeled as “Unknown”; In stark contrast, Science—which comprises Physics and Biology topics—faces significant challenges even when the samples are verifiable.
[…]
FLD (CS) 37.27 10.65 26.62 $\downarrow$
FLD (DS) 18.88 13.64 5.24 $\uparrow$
Arts (CS) 18.60 16.06 2.54 $\uparrow$
Arts (DS) 22.54 14.65 7.89 $\downarrow$
Table 4: Different Stimulating Method Influence on Converting Rate, where $\mathrm{TCR}^{1}=\mathrm{TCR}_{u}^{1}+\mathrm{TCR}_{v}^{1}$. CS, DS refer to Concise and Detailed Stimulation. Rows shaded gray denote the style used in our framework.
Stimulating Style Comparison
For Concise Stimulation, the approach is cost-efficient but may result in less optimal outcomes. On the other hand, Detailed Stimulation can be more effective due to their detailed guidance, though they incur a higher cost because multiple conditions are evaluated step by step. As such, we employed Detailed Stimulation for experiments involving two input conditions, and Concise Stimulation for those with three or more input conditions.
In order to evaluate two styles of prompt stimulation, we swapped their formats and applied them to the corresponding inverse parts of our framework, using GPT-4o as the model.
[…]
Besides, we dive into whether changing stimulation style may lead to aligned distribution changes. That is, as the ability to distinguish Verifiable samples getting weaker, the ability to distinguish Unverifiable samples would get stronger, we denote this is because of the Reasoning Ability change.
Ablation Study
From the Stage 1 output, we isolated the Under-Inspired samples—those that failed to convert—and fed them back through the same Stage-1 prompting protocol.
As Table 4 shows, this second pass produces virtually no additional conversions: more than 95% of the samples remain unchanged. This finding demonstrates that Vague Perception is largely immune to repeated, identical n-shot prompting; applying the same stimulation twice is therefore futile. Consequently, any effective correction must rely on a methodologically distinct Stage 2 rather than reiterating Stage 1.
Phenomenon Exploring
To probe the root causes of GPT-4o’s mistakes, we annotate each error instance with one of four labels:
  1. 1.
    Fact Understanding(FU) – the model does not correctly retrieve, align, or synthesise the relevant facts stated in the passage.
  2. 2.
    Reasoning Gap (RG) – the model retrieves the facts but fails to perform the intermediate logical operations that connect those facts to the hypothesis.
  3. 3.
    Else – a catch-all for infrequent factors (e.g. ignoring task-specific constraints).
Figure 5: Distribution of the root causes behind GPT-4o’s Vague Perception errors across all benchmarks. “FU”: Fact Understanding (failure), “RG”: Reasoning Gap (logical), and “Else” aggregates residual minor factors. A detailed analysis refers to Sec. Phenomenon Exploring Phenomenon Exploring.
Models GPT‑4o GPT‑3.5
Metrics(%) TCR2 Deg TCR2 Deg
FLD 28.19$\uparrow$ 3.08 14.43$\downarrow$ 3.09
FOLIO 44.40$\uparrow$ 5.22 42.10$\downarrow$ 6.32
Science 7.76$\uparrow$ 26.86 3.33$\downarrow$ 3.33
Arts 8.24$\uparrow$ 1.13 4.00$\downarrow$ 0.00
Table 5: Performance comparison of Degrading Analysis Phenomenon between GPT‑4o and GPT‑3.5-turbo-1106.
Quantitative Analysis of Error Drivers.
Figure 5 presents the distribution of the three plotted root causes (FU, RG, Else); which manifests as a distinct abstention behavior (1) FLD and SciQA–Science—are overwhelmingly dominated by FU (87 % and 69 %), with RG contributing only 13 % and 27 %.
In sharp contrast, FOLIO and SciQA–Arts—show near parity between the two main factors: FU slightly lower than RG (40% vs. 51% in FOLIO; 44% vs. 50% in SciQA–Arts).
Across the latter benchmarks, the residual Else category remains minor (2.4-5.4 %).
Interpretation
Benchmarks that demand specialized domain knowledge, such as FLD and SciQA–Science, chiefly test a model’s ability to retrieve and ground domain-specific facts. Benchmarks built on everyday language—such as FOLIO and SciQA–Arts—place a lighter burden on factual retrieval and instead foreground the model’s intrinsic reasoning.
Excessive Caution
Our experiments indicate that Excessive Caution—instances where the model answers “Unknown” despite having enough evidence—is virtually absent in Fact-based benchmarks. However, its incidence, rises from 2.4% in SciQA–Science to 5.4% in SciQA–Arts.
We attribute this increase to the narrative nature of Story-based tasks: unlike Fact-based benchmarks, which provide atomic, prestructured information, the long passages in SciQA–Arts intertwine literary allusions and background knowledge, amplifying epistemic uncertainty. Confronted with such ambiguity, GPT-4o often prefers a hedged response over a potentially erroneous definitive claim, illustrating a trade-off between safety and decisiveness.
Implications.
FLD records the lowest share of reasoning errors. Meanwhile, even though FOLIO exhibit relatively higher RG rate, the total number of samples is far lower than all other benchmarks as shown in Fig 4. These findings indicate that Fact-based benchmarks substantially mitigate Vague Perception by presenting problems as atomic, explicitly structured facts that anchor the model’s reasoning. Compared with story-based tasks, they markedly suppress Excessive Caution and yield fewer instances classified as RG.
Major leaps in raw reasoning competence alone may be limited in the near future. Our findings suggest that even though improving reasoning ability remains the primary bottleneck, we can still ease the problem of Vague Perception by elaborately designed training strategy.
Degrading Analysis
Among the Pipeline, we don’t leave out the possibility that the Previous mislabeled samples, FC2 from Stage 1 may be converted to Unknown, that should be denoted as UC2. As it is reasonable that at first, the model may generate an incorrect answer, but after further stimulating, due to its limited reasoning ability, it fails to arrive at the correct answer and can only output ’Unknown’. We denote this concept as follows:
$Deg=\frac{\left|FC^{1}\setminus\left(FC^{2}\cup TC^{2}\right)\right|}{\left|FC^{1}\right|}$
As shown in Table 5, only a small proportion of samples exhibit this phenomenon. However, in the case of GPT-4o on the reasoning-based Science dataset, the issue becomes notably more pronounced.
Specifically, within the Science dataset, GPT-4o achieves the highest TCR2 among all four datasets. Coupled with a higher Deg value, this suggests that when confronted with problems it previously answered incorrectly—especially those requiring reasoning beyond its capability—GPT-4o tends to respond with an ”Unknown” label. In contrast, GPT-3.5 is more likely to provide a definitive but incorrect answer rather than acknowledge uncertainty.
Conclusion
We introduce WakenLLM, a unified and fine-grained framework for evaluating the Vague Perception phenomenon of Large Language Models. By integrating overall benchmark accuracy, stepwise soundness, and representation-level probing, WakenLLM enables one step further into stimulating LLMs to better reason and expose their potential, thus improving reasoning tasks accuracy. Leveraging this framework, we conduct a systematic investigation of how this phenomenon occurs and how to inspire the model to a larger extent.
Our experiments demonstrate that although models may not arrive at the correct answer on their first attempt, targeted stimulation can reliably guide them toward improved reasoning and significantly higher final accuracy. Furthermore, conformity experiments reveals that the model’s ability to predict Unknown labels can be much easier to mis-stimulate than T/F labels, influencing the result. These findings offer practical guidance for designing training strategies to ease the problem of vague answering and finally provide robust answers for daily use cases. Ultimately, our work highlights the importance of evaluating LLM reasoning systems not only on their ”one-time” performance but also on their capacity for ”stimulated” self-correction.